# Physics-aware Machine Learning Revolutionizes Scientific Paradigm for Machine Learning and Process-based Hydrology

### v1

@@ 2.1.2 Deep Operator Networks
Operator learning strategies are predicated on the principles of the universal approximation theorem [76]. This approximation ability has formed the basis of modern deep operator network development. Here, we present an introduction to deep operator networks (DeepONet) [76], the Fourier neural operator (FNOs) [77], the wavelet neural operators (WNOs) [78], and the graph neural operators (GNOs) [79].
@@
DeepONets utilize the aforementioned branch-trunk architecture in order to map finite inputs to the response space [76]. This enables domain encoding to be handled by the trunk network, with any number of branch networks used to handle the various inputs associated with the problem. For example, in [80], knowledge of the dynamics of decaying evolution of the Brusselator reaction-diffusion system is incorporated by employing a trigonometric feature expansion of the temporal input of the trunk network. In addition, to employ DeepONet for realistic setups (multiple input functions) and diverse applications, a multiple input DeepONet [81] is presented for multiple initial conditions and boundary conditions at the same time. In addition, to enable the learning of the DeepONet with a smaller set of parameters, HyperDeepONet [82] is proposed by using the expressive power of the hyper network. However, several studies point out that the slow decay rate of the lower bound leads to inaccurate approximation operator learning for complex physical dynamics using DeepONet [83, 84].
@@
FNOs [77, 85] use Fourier Transformations to move to the infinite-dimensional response space. The operator takes input functions specified on a well-defined equally spaced lattice grid and outputs the field of interest on the same grid points. The network parameters are defined and learned in the Fourier space rather than in the physical space. FNO has become one of the mainstream physical data-guided deep operator networks used to solve PDEs [38]. Specifically, to solve PDEs on arbitrary geometries, Geo-FNO [86] is designed by learning to deform the input (physical) domain. Domain agnostic FNO [87] is proposed by incorporating a smoothed characteristic function in the integral layer architecture of FNOs. However, the FNO and the geo-FNO perform worse and are unstable on complex geometries and noisy data. Thus, factorized FNO [88] is proposed to improve representation stability for the operator. In addition, a host of variants of FNOs are proposed to improve the generalization of FNOs for solving PDEs, such as incremental FNO [89].
@@
One major shortcoming of FNOs is that the basis functions in FFTs are generally frequency localized with no spatial resolution [90]. Thus, WNOs are proposed, to learn the network parameters in the wavelet space that are both frequency and spatial localized, and thereby learn the signal patterns more effectively. Recently, a coupled multiwavelets neural operator learning scheme [91] is proposed to solve coupled PDEs, by decoupling the coupled integral kernels in the wavelet space. This approach has shown that WNO can handle domains with both smooth and complex geometries, and has been applied in learning solution operators for a highly nonlinear family of PDEs with discontinuities and abrupt changes in the solution domain and boundary.
@@
GNOs [79] are computed by message passing on graph networks, which are able to capture the graph (non-local) structure of the data. Graph kernel networks (GKNs) are mainly used in GNOs. The idea of GKNs comes from parameterizing Green’s function in an iterative architecture [79]. As an integral neural network similar to FNOs, GKNs are composed of a lifting layer, iterative kernel integration layers, and a projection layer. However, it has been found that GKNs might be unstable when increasing the number of kernel integration layers [92]. Thus, a nonlocal neural operator [92] is proposed by using a resolution-independent nonlocal kernel network.
@@ 2.1.3 Physics-discovery Neural Networks
The discovery of dynamic behavior is crucial, as it enables us to comprehend the underlying physical laws that regulate complex systems. However, discovering governing equations is a challenging task due to complex and nonlinear systems, as well as noisy and incomplete data. With the rapid advance of machine intelligence [12], the past three decades have seen a considerable amount of physics-discovery neural networks growing up around the interplay between ML and scientific discovery. Essentially, two genres have been observed based on unavailable and available prior knowledge of the physics system. First, the natural and optimal solution to automatically identifying the governing equations for dynamic systems is to learn a symbolic model from experimental data. Hence, symbolic regression [93] and symbolic neural networks (SNNs) [94, 95, 96] have been explored for inferring concise equations without any prior knowledge. A physics-inspired method called AI Feynman [97, 98] is proposed for symbolic regression. AI Feynman employs NNs to discover generalized symmetries in complex data, enabling the recursive decomposition of difficult problems into simpler ones with fewer variables. GNN [99] is also utilized to identify the nontrivial relation due to its excellent inductive capability. In addition, PDgML can apply for knowledge discovery that cannot be modeled from a physical process-based perspective, such as the two-way feedback between human and water systems [6, 7]. Second, discovering dynamic behavior from data is often accomplished by defining a large set of possible mathematical basis functions or process-based models based on prior knowledge. The most representative works are SINDy [100] and PDE-FIND [101] for ordinary differential equations (ODEs) and PDEs, respectively. However, the standard sparse representation-based methods are limited to high-fidelity noiseless measurement, which is usually difficult and expensive to obtain. Hence, many recent efforts have been devoted to discovering PDEs from sparse/noisy data [102, 103].
Table 1: Proposed Physical Data-guided Machine Learning
Categories Types Examples
Objective(s)
Limitation(s)
Classical Physical Data-guided Neural Networks Spatio-temporal modeling LSTM [42]
Three gates to keep long-standing relevant information and discard irrelevant information
Limited generalization, high computational cost, limited physical interpretability
Physics-guided LSTM [43]
A deep learning network for rainfall-runoff simulation
Informer [45]
Long sequence time-series forecasting
ConvLSTM [49]
A spatio-temporal sequence forecasting problem
Earthformer [53]
A space-time Transformer for Earth system forecasting
Geometric deep learning [57]
A spatially-irregular forecast method
Dynamical system modeling PDE-Net [60]
A feed-forward deep network-based PDE solver
Unable to handle temporally or spatially sparse or non-uniform observations
Mesh-Conv [65]
A solver for modeling non-uniform structured or unstructured spatial mesh
High computational cost
Neural message passing-based models [66, 67, 68, 69, 70, 71, 72]
A general framework for supervised learning on graphs (arbitrary domains)
Mesh-based or particle-based; message-passing between small-scale moving and interacting objects
DRL-based method [75]
A coupled 2D control system
Limited capabilities and robustness; unknown behavior and convergence speed in action spaces of high dimensionalities
Deep Operator Networks Deep Operator Networks (DeepONet) [76] Multiple input DeepONet [81]
Employ DeepONet for realistic setups (multiple initial conditions and boundary conditions)
Inaccurate approximation for complex physical dynamics
HyperDeepONet [82]
Enable the learning of the DeepONet with a smaller set of parameters
Fourier neural operator (FNOs) [77] Geo-FNO [86], Domain agnostic FNO [87]
Solve PDEs on arbitrary geometries
Unstable, limited generalization, the basis functions in FFTs are generally frequency localised with no spatial resolution
Factorized FNO [88]
Improve representation stability for the operator
Wavelet neural operatos (WNOs) [78] WNO [78]
Learn the network parameters in the wavelet space
Complexity, limited interpretability
Coupled multiwavelets neural operator [91]
Learn coupled PDEs in the wavelet space
Graph neural operators (GNOs) [79] Graph kernel networks [79]
An integral neural network that parameterizes Green’s function
Difficult to train, computational cost, limited generalization
Nonlocal neural operator [92]
A resolution-independent nonlocal kernel network
Physics-discovery Neural Networks Physics-discovery with unavailable prior knowledge Symbolic neural networks (SNNs) [94, 95, 96]
Infer concise equations without any prior knowledge
Challenging task due to complex and nonlinear systems, as well as noisy and incomplete data
AI Feynman [97, 98]
Discover generalized symmetries in complex data
GNN [99]
Identify the nontrivial relation due to its excellent inductive capability
Knowledge-agnostic discovery [6, 7]
Discover the two-way feedback between human and water systems
Physics-discovery with available prior knowledge SINDy [100]
Discover governing equations from data by sparse identification of nonlinear dynamical systems
Limited high-fidelity noiseless measurement
PDE-FIND [101]
Data-driven discovery for PDEs
@@ 2.2 Physics-informed Machine Learning
In scientific computing, physical phenomena are often described using a strong mathematical form consisting of governing differential equations as well as initial and boundary conditions. PiML is a widely-used approach that integrates physical laws into ML models (for example, designing loss functions or regularization), facilitates the accurate capture of dynamic patterns and concomitantly diminishes the search space for model parameters. This approach is sometimes referred to as imposing differentiable constraints in loss functions. It integrates (noisy) data and mathematical models, and implements them through neural networks (physics-informed neural networks) or kernel-based neural operators (physics-informed neural operators). The objectives and limitations of different physics-informed machine learning are briefly discussed in Table 2.2.
@@
Categories
Types
Examples
Objective(s)
Limitation(s)
Vanilla PINNs
PINNs [25]
A deep learning framework for solving forward and inverse problems involving nonlinear partial differential equations
@@
The challenging optimization landscape from soft physics or PDE constraints; the difficulty of propagating information from the initial or boundary conditions to unseen parts of the interior or to future times; the sensitivity to hyperparameter selection; the slow training/re-optimization process in practical applications and different instances; failing to converge on high-frequency, multi-scale PDEs, or time-dependent PDEs
Problem-specific insights
Universal PINNs [104]
@@
Solve PDEs with a point source by a lower bound constrained uncertainty weighting algorithm and a multi-scale deep neural network
Limited accuracy and generalization
cPINN [105]
@@
Develop a unique conservative physics-informed neural network on discrete domains for nonlinear conservation laws
Limited representation and parallelization capacity
nPINNs [106]
@@
Extend PINNs to parameter and function inference for integral equations such as nonlocal Poisson and nonlocal turbulence models
Challenging optimization; require more residual points
Improved PINNs [107]
@@
An improved PINN approach for localized wave solutions of the derivative nonlinear Schrodinger equation in complex space with faster convergence and optimum simulation performance
Lack of consideration of complex integrable equations
AdaptivePINNs [108]
@@
An automatic numerical solver for the Allen-Cahn and Cahn-Hilliard equations that proposes an improved PINN
Limited application range (focused on the problem of solving differential equations)
CINN [109]
@@
A simple and efficient characteristic-informed neural networks for solving forward and inverse problems involving hyperbolic PDEs
Limited application; limited expression and generalization ability
NSFnets [110]
Physics-informed neural networks for the incompressible Navier-Stokes equations by automatic differentiation
@@
Limited expressivity and generalization; hyperparameters; failure to converge on high-frequency PDEs; high computational cost
GAPINN [111]
@@
Geometry aware physics-informed neural network for solving Navier–Stokes equation on irregular geometries
Limited expressivity of framework; cannot deal with time varying geometries
CPINNs [112]
@@
Coupled physics-informed neural networks for the closed-loop geothermal system
Limited expressivity of framework; hyperparameters
Input data
Bayesian-PINNs [113]
@@
Solve both forward and inverse nonlinear problems described by PDEs and noisy data
Limited representation; lack of large scale dataset testing
ModalPINNs [114]
@@
An extension of physics-informed neural networks with enforced truncated Fourier decomposition for periodic flow reconstruction using a limited number of imperfect sensors
Limited application range
$\Delta$-PINNs [115]
A positional encoding mechanism for PINNs based on the eigenfunctions of the Laplace-Beltrami operator, presented on complex geometries
@@
Limited accuracy using finite elements; computing eigenfunctions of the Laplace-Beltrami operator can become expensive for large meshes
Frameworks
PhyCRNet [116]
@@
An encoder-decoder, ConvLSTM, is proposed for low-dimensional spatial feature extraction and temporal evolution learning. The loss function is defined as the aggregated discretized PDE residuals
Limited accuracy due to finite difference
PINNs with ensemble models [117]
Train an ensemble of PINNs and use the ensemble agreement (confidence) as the criterion for expanding the solution interval to new areas
High computational cost
Physics-informed GNNs [118]
@@
Physics-informed graph neural networks not only provide a representation that is specially tailored to the symmetries and properties of the systems we consider, but they also are naturally invariant to the extent of the system
High computational cost; limited to homogeneous systems
MH-PINNs [119]
@@
Multi-head physics-informed neural network constructed as a potent tool for multi-task learning, generative modeling, and few-shot learning for diverse problems
Limited expressivity of framework; hyperparameters; difficult to optimize
Phase2vec [120]
@@
An embedding method that learns high-quality, physically-meaningful representations of 2D dynamical systems by a convolutional backbone
Failure in the n-dimensional case; dynamics are qualitatively invariant to translations of the input coordinate system
PI-PINN [121]
@@
Orthogonal polynomials are used to construct the neural network for solving nonlinear partial differential equations
Limited expressivity of framework
PhyGeoNet [122]
@@
Physics-constrained CNN learning architecture to learn solutions of parametric PDEs on irregular domains without any labeled data.
Limited accuracy due to finite difference; cannot solve dynamic geometrically irregular domains
Spline-PINN [123]
@@
Approaching PDEs training without any precomputed training data using a physics-informed loss function only; provides fast, continuous solutions that generalize to unseen domains
Limited accuracy due to finite difference
Physics-constrained Unsupervised Learning [124]
@@
Represent meshes naturally as graphs, process them using graph networks, and formulate physics-based loss to provide an unsupervised learning framework for PDEs
Limited accuracy due to finite difference
Improving speed and efficiency
Meta-learning based PINNs
@@
de Avila Belbute-Peres et al. [125] meta-trains a hyper-network that for each task can generate weights for a small neural network. Psaros et al. [126] meta-learns a loss function that is used to optimize the NN. Penwarden et al. [127] proposes a meta-learning approach that learns an initialization of weights such that the NN can be optimized quickly.
Limited expressivity of framework, unsteady in solving related PDEs
XPINNs [128]
@@
A generalized space-time domain decomposition framework, proposed for PINNs to solve nonlinear PDEs on arbitrary complex-geometry domains
Hyperparameters; difficult to optimize
canPINN [129]
@@
Coupled-automatic-numerical differentiation framework that unifies the advantages of automatic differentiation and numerical differentiation, providing more robust and efficient training than automatic differentiation-based PINNs
Limited application scenarios (generally applicable to non-uniform samplings), limited computational efficiency
DT-PINNs [130]
@@
DT-PINNs are trained by replacing these exact spatial derivatives with high-order accurate numerical discretizations computed using meshless radial basis function-finite differences and applied via sparse-matrix vector multiplication
Limited accuracy due to radial basis function-finite differences
Improving generalization
TPINN [131]
@@
Transfer physics-informed neural network for solving forward and inverse problems in nonlinear partial differential equations by parameter sharing
Lack of consideration noise in input data
Multiscale and multiphysics problems
FBPINNs [132]
@@
Finite basis physics-informed neural networks use ideas from domain decomposition to accelerate the learning process of PINNs and improve their accuracy
Fail to coverge on high-frequency PDEs; limited accuracy and expressivity
Curriculum and Sequence-to-sequence PINN [133]
@@
Curriculum regularization and sequence-to-sequence learning, used to address the problems of PINNs that fail to learn relevant physical phenomena for even slightly more complex problems
Limited expressivity of framework; hyperparameters; lack of realistic complex application
Multi-scale PINNs [134]
@@
Spatio-temporal and multi-scale random Fourier features are used to address high-frequency or multi-scale problems in PINN
Hyperparameters
Hyperparameter selection
Self-adaptation PINNs [135]
@@
Self-adaptation PINNs that make weights increase as the corresponding losses increase, which is accomplished by training the network to simultaneously minimize losses and maximize weights
Limited expressivity of framework; the squared residual penalty method imposes a fundamental limitation associated with conditioning
Heuristic method [136]
@@
A heuristic method for optimizing weighted loss functions for solving PDEs with NNs
Data-physics-driven parameter discovery
PINN-TFC [137]
@@
Data-physics-driven parameters discovery of problems modeled via ordinary differential equations
Cannot be applied to the data-driven discovery of problems
PINN-SR
@@
Physics-informed neural network with sparse regression to discover governing partial differential equations from scarce and noisy data for nonlinear spatio-temporal systems
High computational cost need a vast number of collocation points for complex systems; limitations of PINNs
Physics-informed Neural Networks
Discovering governing equations [138]
@@
Solve the inverse problem of parameter identification in discrete, high-dimensional systems inspired by physical applications
Limitations of PINNs; cannot address higher dimensional settings
DeepONet-based
Physics-informed DeepONet [139]
@@
Train the DeepONet by directly incorporating known differential equations into the loss function, along with some labeled datasets
Limitations on fixed grid or basis; cannot achieve discretization invariant; limitation of linear approximation; limited computational efficiency and generalization
FNO-based
PINO [140]
An integration of Fourier operator learning and physics-informed settings
@@
Limited accuracy on solving non-periodic problems; limited adaptability due to the selection of model parameters; it cannot be applied to large-scale and long-term sequence PDE solution
FC-PINO [141]
An architecture that leverages Fourier continuation is proposed to apply the exact gradient method to PINO for nonperiodic problems
@@
Limited adaptability due to the selection of model parameters; it cannot be applied to large-scale and long-term sequence PDE solution; it cannot achieve geometry-adaptive learning
GeoPINS [142]
@@
A geometry-adaptive physics-informed neural solver based on the advantages of no training data in physics-informed neural networks, as well as possessing a fast, accurate, geometry-adaptive and resolution-invariant architecture
Limited adaptability due to the selection of model parameters
Physics- informed Neural Operators
WNO-based
Physics-informed WNO [143]
A physics-informed WNO for learning the solution operators of families of parametric PDEs without labeled training data, based on the advantage of time-frequency localization of wavelets
@@
Complexity; the absence of a standardized procedure to obtain the optimum and desired network architecture of a physics-informed WNO and to assign the weights of boundary/initial condition loss; the lack of information regarding how to learn the operator for systems involving multi-scale physics systems

### v4

@@ 2.1.2 Deep Operator Networks
Operator learning strategies are based on the principles of the universal approximation theorem [47]. This capability to approximate functions forms the foundation of modern developments in deep operator networks. Here, we present an introduction to deep operator networks (DeepONet) [47], the Fourier neural operator (FNOs) [50], the wavelet neural operators (WNOs) [54], and the graph neural operators (GNOs) [56].
@@
DeepONets employ the aforementioned branch-trunk architecture to map finite inputs to the response space [47]. This allows the trunk network to encode the domain, with multiple branch networks handling various inputs relevant to the problem. For instance, in [95], the decaying dynamics of the Brusselator reaction-diffusion system are incorporated by utilizing a trigonometric feature expansion of the temporal input in the trunk network. Additionally, to enable the application of DeepONet in realistic scenarios involving multiple input functions and diverse applications, a multiple input DeepONet [48] is introduced to handle multiple initial conditions and boundary conditions simultaneously. In addition, to enable the learning of the DeepONet with a smaller set of parameters, HyperDeepONet [49] is proposed by using the expressive power of the hyper network. However, several studies point out that the slow decay rate of the lower bound leads to inaccurate approximation operator learning for complex physical dynamics using DeepONet [96, 97].
@@
FNOs [50, 98] use Fourier Transformations to move to the infinite-dimensional response space. These operators process input functions defined on a well-defined, equally spaced lattice grid and produce the desired fields at the same grid points. The network parameters are established and trained in the Fourier domain instead of the physical space. FNO has become one of the mainstream physical data-guided deep operator networks used to solve PDEs [73]. Specifically, to solve PDEs on arbitrary geometries, Geo-FNO [51] is designed by learning to deform the input (physical) domain. Domain agnostic FNO [52] is proposed by incorporating a smoothed characteristic function within the integral layer architecture of FNOs. However, the FNO and the geo-FNO perform worse and are unstable on complex geometries and noisy data. Thus, factorized FNO [53] is proposed to improve representation stability for the operator. In addition, a host of variants of FNOs are proposed to improve the generalization of FNOs for solving PDEs, such as incremental FNO [99].
@@
A significant limitation of FNOs is that FFT basis functions are typically frequency localized without spatial resolution [100]. Thus, WNOs are proposed, to learn network parameters in the wavelet space, which are both frequency and spatially localized, thereby enabling more effective learning of signal patterns. Recently, a coupled multiwavelets neural operator learning scheme [55] is proposed to solve coupled PDEs, by decoupling the coupled integral kernels in the wavelet space. This approach has demonstrated that WNO can effectively manage domains with both smooth and complex geometries. It has been utilized to learn solution operators for a highly nonlinear family of PDEs characterized by discontinuities and abrupt changes in both the solution domain and its boundary.
@@
GNOs [56] utilize message passing on graph networks to capture the non-local structure of data. Graph kernel networks (GKNs), which originate from parameterizing Green’s functions in an iterative architecture [56], are primarily employed in GNOs. Similar to FNOs, GKNs consist of a lifting layer, iterative kernel integration layers, and a projection layer. However, recent studies have noted that GKNs may exhibit instability with an increased number of iterative kernel integration layers [57]. Therefore, a resolution-independent nonlocal neural operator [57] has been proposed to address these issues.
@@ 2.1.3 Physics-discovery Neural Networks
Discovering dynamical processes is crucial as it enables us to uncover the fundamental physical laws that govern complex systems. However, the discovery of the governing equations presents a significant challenge due to the inherent complexity and nonlinearity of these systems, as well as the presence of noisy and incomplete data. With the rapid advancement of machine intelligence [12], the past three decades have witnessed significant development in physics-discovery neural networks, at the intersection of ML and scientific discovery. Two main approaches have emerged based on unavailable and available prior knowledge of the physics system. First, the natural and optimal solution to automatically identifying the governing equations for dynamic systems is to learn a symbolic model from experimental data. Symbolic regression [101] and symbolic neural networks (SNNs) [58, 59, 60] have been explored extensively for inferring concise equations without prior knowledge. A physics-inspired method called AI Feynman [61, 62] is proposed for symbolic regression. AI Feynman employs NNs to discover generalized symmetries in complex data, enabling the recursive decomposition of difficult problems into simpler ones with fewer variables. GNN [63] is also utilized to identify the nontrivial relation due to its exceptional inductive capability. In addition, PDgML is applied for knowledge discovery that cannot be modeled from a physical process-based perspective, such as the two-way feedback between human and water systems [6, 7]. Second, discovering dynamic behavior from data often involves defining a large set of possible mathematical basis functions or process-based models based on prior knowledge. Representative works include SINDy [64] and PDE-FIND [65] for ordinary differential equations (ODEs) and PDEs, respectively. However, these standard sparse representation-based methods are typically limited to high-fidelity noiseless measurements, which are often difficult and expensive to obtain. Consequently, recent efforts have focused on discovering PDEs from sparse or noisy data [102, 103].
@@ 2.2 Physics-informed Machine Learning
In scientific computing, physical phenomena are typically described using a robust mathematical framework that includes governing differential equations, as well as initial and boundary conditions. PiML is a widely-used approach that integrates physical laws into ML models (for example, designing loss functions or regularization), facilitates the accurate capture of dynamic patterns and concomitantly diminishes the search space for model parameters. This approach is sometimes referred to as imposing differentiable constraints in loss functions. It integrates (noisy) data and mathematical models, and implements them through neural networks (physics-informed neural networks) or kernel-based neural operators (physics-informed neural operators). The objectives and limitations of different physics-informed machine learning are briefly discussed in Table 2.2.
@@
Categories Types Examples Objective(s) Limitation(s)
@@
Vanilla PINNs PINNs [25] A deep learning framework designed to address forward and inverse problems associated with nonlinear partial differential equations The challenging optimizations due to soft physics or PDE constraints; the challenge of nonlinear partial differential equations conveying information from initial or boundary conditions to unseen interior regions or future times; the sensitivity to hyperparameter selection; the slow training and re-optimization processes in practical applications and various instances; and the failure to converge on high-frequency, multi-scale, or time-dependent PDEs
@@
Problem-specific insights Universal PINNs [104] Solve PDEs with a point source by a lower bound constrained uncertainty weighting algorithm and a multi-scale deep neural network Limited accuracy and generalization
@@
cPINN [105] Create a distinct conservative physics-informed neural network tailored for nonlinear conservation laws on discrete domains Limited representation and parallelization capacity
@@
nPINNs [106] Extend PINNs to parameter and function inference for integral equations such as nonlocal Poisson and nonlocal turbulence models Challenging optimization; require more residual points
@@
Improved PINNs [107] An improved PINN method for localized wave solutions of the derivative nonlinear Schrödinger equation in complex space, offering quicker convergence and optimal simulation performance Lack of consideration of complex integrable equations
@@
AdaptivePINNs [108] An automatic numerical solver for the Allen-Cahn and Cahn-Hilliard equations that utilizes an improved PINN Limited application range (focused on the problem of solving differential equations)
@@
CINN [109] A straightforward and efficient characteristic-informed neural networks for solving forward and inverse problems in hyperbolic PDEs Limited application; limited expression and generalization ability
@@
NSFnets [110] Physics-informed neural networks for the incompressible Navier-Stokes equations by automatic differentiation Limited expressivity and generalization; hyperparameters; failure to converge on high-frequency PDEs; high computational cost
@@
GAPINN [111] Geometry aware physics-informed neural network for solving Navier–Stokes equation on irregular geometries Limited expressivity of framework; cannot deal with time varying geometries
@@
CPINNs [112] Coupled physics-informed neural networks designed for closed-loop geothermal systems Limited expressivity of framework; hyperparameters
@@
Input data Bayesian-PINNs [113] Solve both forward and inverse nonlinear problems described by PDEs and noisy data Limited representation; lack of large scale dataset testing
@@
ModalPINNs [114] An extension of physics-informed neural networks incorporating enforced truncated Fourier for periodic flow reconstruction with a limited number of imperfect sensors Limited application range
@@
$\Delta$-PINNs [115] A positional encoding mechanism for PINNs based on the eigenfunctions of the Laplace-Beltrami operator, presented on complex geometries Limited accuracy using finite elements; computing eigenfunctions of the Laplace-Beltrami operator can become expensive for large meshes
@@
Frameworks PhyCRNet [116] An encoder-decoder, ConvLSTM, is proposed for low-dimensional spatial feature extraction and temporal evolution learning. The loss function is defined as the aggregated discretized PDE residuals Limited accuracy due to finite difference
PINNs with ensemble models [117] Train a set of ensemble PINNs and utilize the consensus (confidence) among the ensemble members as the criterion for extending the solution range into new domains High computational cost
@@
Physics-informed GNNs [118] Physics-informed graph neural networks not only provide a representation that is specially tailored to the symmetries and properties of the systems we consider, but they also are naturally invariant to the extent of the system High computational cost; limited to homogeneous systems
@@
MH-PINNs [119] Multi-head physics-informed neural network constructed as a potent tool for multi-task learning, generative modeling, and few-shot learning for diverse problems Limited expressivity of framework; hyperparameters; difficult to optimize
@@
Phase2vec [120] An embedding method that learns high-quality, physically-meaningful representations of 2D dynamical systems by a convolutional backbone Failure in the n-dimensional case; dynamics are qualitatively invariant to translations of the input coordinate system
@@
PI-PINN [121] Orthogonal polynomials are used to construct the neural network for solving nonlinear partial differential equations Limited expressivity of framework
@@
PhyGeoNet [122] Physics-constrained CNN learning architecture to learn solutions of parametric PDEs on irregular domains without any labeled data. Limited accuracy due to finite difference; cannot solve dynamic geometrically irregular domains
@@
Spline-PINN [123] Training PDEs without any precomputed data using only a physics-informed loss function; provides rapid, continuous solutions that generalize across unseen domains Limited accuracy due to finite difference
@@
Physics-constrained Unsupervised Learning [124] Represent meshes naturally as graphs, process them using graph networks, and formulate physics-based loss to provide an unsupervised learning framework for PDEs Limited accuracy due to finite difference
@@
Improving speed and efficiency Meta-learning based PINNs de Avila Belbute-Peres et al. [125] meta-trains a hyper-network that for each task can generate weights for a small neural network. Psaros et al. [126] meta-learns a loss function that is used to optimize the NN. Penwarden et al. [127] proposes a meta-learning approach that learns an initialization of weights such that the NN can be optimized quickly. Limited expressivity of framework, unsteady in solving related PDEs
@@
XPINNs [128] A generalized space-time domain decomposition framework, proposed for PINNs to solve nonlinear PDEs on arbitrary complex-geometry domains Hyperparameters; difficult to optimize
@@
canPINN [129] Coupled-automatic-numerical differentiation framework that unifies the advantages of automatic differentiation and numerical differentiation, providing more robust and efficient training than automatic differentiation-based PINNs Limited application scenarios (generally applicable to non-uniform samplings), limited computational efficiency
@@
DT-PINNs [130] DT-PINNs are trained by substituting precise spatial derivatives with high-order accurate numerical discretizations calculated using meshless radial basis function-finite differences and implemented via sparse matrix-vector multiplication Limited accuracy due to radial basis function-finite differences
@@
Improving generalization TPINN [131] Transfer physics-informed neural network for solving forward and inverse problems in nonlinear partial differential equations by parameter sharing Lack of consideration noise in input data
@@
Multiscale and multiphysics problems FBPINNs [132] Finite basis physics-informed neural networks use ideas from domain decomposition to accelerate the learning process of PINNs and improve their accuracy Fail to coverge on high-frequency PDEs; limited accuracy and expressivity
@@
Curriculum and Sequence-to-sequence PINN [133] Curriculum regularization and sequence-to-sequence learning, used to address the problems of PINNs that fail to learn relevant physical phenomena for even slightly more complex problems Limited expressivity of framework; hyperparameters; lack of realistic complex application
@@
Multi-scale PINNs [134] Spatio-temporal and multi-scale random Fourier features are used to address high-frequency or multi-scale problems in PINN Hyperparameters
@@
Hyperparameter selection Self-adaptation PINNs [135] Self-adaptation PINNs that make weights increase as the corresponding losses increase, which is accomplished by training the network to simultaneously minimize losses and maximize weights Limited expressivity of framework; the squared residual penalty method imposes a fundamental limitation associated with conditioning
@@
Heuristic method [136] A heuristic method for optimizing weighted loss functions for solving PDEs with NNs
@@
Data-physics-driven parameter discovery PINN-TFC [137] Data-physics-driven parameters discovery of problems modeled via ordinary differential equations Cannot be applied to the data-driven discovery of problems
@@
PINN-SR Physics-informed neural network with sparse regression to discover governing partial differential equations from scarce and noisy data for nonlinear spatio-temporal systems High computational cost need a vast number of collocation points for complex systems; limitations of PINNs
@@
Physics-informed Neural Networks Discovering governing equations [138] Solve the inverse problem of parameter identification in discrete, high-dimensional systems inspired by physical applications Limitations of PINNs; cannot address higher dimensional settings
@@
DeepONet-based Physics-informed DeepONet [139] Train the DeepONet by directly incorporating known differential equations into the loss function, along with some labeled datasets Limitations on fixed grid or basis; cannot achieve discretization invariant; limitation of linear approximation; limited computational efficiency and generalization
@@
FNO-based PINO [140] An integration of Fourier operator learning and physics-informed settings Limited accuracy on solving non-periodic problems; limited adaptability due to the selection of model parameters; it cannot be applied to large-scale and long-term sequence PDE solution
@@
FC-PINO [141] An architecture that leverages Fourier continuation is proposed to apply the exact gradient method to PINO for nonperiodic problems Limited adaptability due to the selection of model parameters; it cannot be applied to large-scale and long-term sequence PDE solution; it cannot achieve geometry-adaptive learning
@@
GeoPINS [142] A geometry-adaptive physics-informed neural solver based on the advantages of no training data in physics-informed neural networks, as well as possessing a fast, accurate, geometry-adaptive and resolution-invariant architecture Limited adaptability due to the selection of model parameters
@@
Physics- informed Neural Operators WNO-based Physics-informed WNO [143] A physics-informed WNO for learning solution operators of families of parametric PDEs without labeled training data, leveraging the time-frequency localization advantage of wavelets. Complexity; the absence of a standardized procedure to obtain the optimum and desired network architecture of a physics-informed WNO and to assign the weights of boundary/initial condition loss; the lack of information regarding how to learn the operator for systems involving multi-scale physics systems